OpenStudy (anonymous):

Please help Evaluate the polynomial function for the given values of the variable P(s)=2sᶟ-9s-10 P(1/2) and P(0)

OpenStudy (anonymous):

p(1/2)= 2 * (1/2)^3 - 9*(1/2) - 10

OpenStudy (anonymous):

P(1/2) = P(s) with s = 1/2

OpenStudy (anonymous):

P(1/2) = -57/4 P(0) = -10

OpenStudy (anonymous):

so set s = 1/2 and 0 in your equation and evaluate it.

OpenStudy (anonymous):

thank you but i would like to work the problem myself step by step so i can do it on my own

OpenStudy (anonymous):

Then just plug in the values 1/2 and 0 to the functions.

OpenStudy (anonymous):

Ok.... do it by your self, and then check the answers using wolfram

OpenStudy (anonymous):

i dont know how i havent done this kind of math in along time will u explain it step by step

OpenStudy (anonymous):

So are you having problems multiplying, adding and subtracting with a fraction?

OpenStudy (anonymous):

yes i am i hate fractions never been good at them

OpenStudy (anonymous):

okay. lets take this one step at a time P(s)=2sᶟ-9s-10 s = 1/2 P(1/2) = 2 * (1/2)^3 - 9(1/2) -10

OpenStudy (anonymous):

(1/2)^3 = 1/2*1/2*1/2 = 1/8 2(1/2)^3 = 2*1/8 = 1/4 9(1/2) = 9/2 so, P(1/2) = 1/4-9/2-10

OpenStudy (anonymous):

9/2 is nothing but 18/4 and 10 is nothing but 40/4 so, P(1/2) = 1/4 - 18/4 - 40/4 = (1-18-40)/4 = -57/4

OpenStudy (anonymous):

2(1/2)^3= how did u get 2*1/8

OpenStudy (anonymous):

what \[(1/2)^3\] its (1/2) *(1/2) *(1/2)= 1/8 but dont forget 2 2/8 = 1/4